OpenStudy (anonymous):

2a+3b=-1, 3a+5b=-2 using the elimination method. please help

OpenStudy (anonymous):

With the elimination method, we want to eliminate one of the variables; 2a + 3b = -1 3a + 5b = -2 We will multiply the first equation with -3 and the second with 2, so we can eliminate the a; -3(2a+3b=-1) 2(3a+5b=-2) new equations; -6a-9b=3 6a+10b=-4 Add the two equations together, and you can eliminate the a; b=-1 Now plug the -1 into one of the original equations; 2a+3(-1)=-1 2a-3=-1 2a=2 a=1 :)

OpenStudy (anonymous):

thank you so much for your time, this is very helpful